A manufacturer knows that their items have a normally distributed lifespan, with a mean of 12.4 years, and standard deviation of 0.6 years.
The 2% of items with the shortest lifespan will last less than how many years?
Give your answer to one decimal place.

Answer by Boreal(15235)

z(0.02)=-2..05
z=(x-mean(/sd)
-2.05=(x-12.4)/0.6
-1.23=x-12.4
x=11.17 years or 11.2 years